150 ml of 0.10 M calcium bromide is added to 150 ml of 0.10 potassium phosphate. what is the the theoretical yield in moles?
1
Expert's answer
2020-02-25T04:53:57-0500

3CaBr2+2K3PO4=Ca3(PO4)2+6KBr

0.015 mol (0.10 M*0.150L) both CaBr2 and K3PO4

Thus, CaBr2 is the limiting reactant

0.015/3= 0.005 mol of Ca3(PO4)2 foms

The theoretical yield is 0.005 mol